# ConCap: Practical Network Traffic Generation for Flow-based Intrusion Detection Systems

### v1

@@ I Introduction
There is an anomaly in research on Network Intrusion Detection Systems (NIDS). On the one hand, a deluge of papers continues to expand the boundaries of our knowledge, as shown by [1, 2, 3]. On the other hand, a growing number of recent efforts highlight some “pitfalls” that undermine the foundations of prior research (e.g., [4, 5, 6, 7, 8]).
@@
To portray this contrast, we can look at the panorama of open-source datasets used in NIDS-related research, some of which count thousands of citations according to Google Scholar [8]. As a concrete example, consider the paper presenting the CICIDS17 dataset [9]: published in 2018, this paper had 1600 citations in Q3 2022, which increased to 4200 in Q4 2024—indicating a substantial growth in NIDS research. Yet, in 2021, Engelen et al. [5] pinpointed glaring issues in CICIDS17—particularly in terms of ground-truth labeling of attack samples. The flaws of CICIDS17 (and also of its successor, CICIDS18) have been “fixed” in 2022 [6]. However, the EuroS&P’24 Best Paper Award [8] revealed that most datasets used by prior research have “bad design smells.”
@@
Simply put, the stark reality is that (i) data is required to support a paper’s claims, but (ii) high-quality data is hard to come by in the NIDS context—especially from the viewpoint of an academic researcher [2, 8, 10]. To aggravate this problem, modern NIDS increasingly rely on data-driven techniques, such as machine learning (ML). Therefore, labeled data is necessary to properly evaluate the pros and cons of state-of-the-art NIDS [2]. Finally, despite the benefits that open-source datasets (under the assumption that they are correctly labeled) can provide to a researcher [11], exclusive reliance on such “static benchmarks” prevents one from generating new data. This impairs the assessment of existing NIDS against recent, and more sophisticated threats—such as multi-step attacks [12, 13], or attacks exploiting recently-discovered vulnerabilities [14]. We aim to rectify such a “data problem.”
@@
Our major technical contribution is ConCap, an open-source system to generate (and label) network-traffic data mimicking that of a real-world network. ConCap is particularly suited to generate network data pertaining to malicious activities. Such data can then be used alongside “benign” data taken from the network environment that the NIDS is meant to protect—which is a well-founded assumption [15, 16, 10]. Practically, such environment can be: (a) that of a benchmark dataset [11], or of (b) an ad-hoc network testbed for experimental research [17], or even that of (c) a real-world network [18]. We design ConCap so that it is (i) flexible enough to enable reproduction of any (allegedly malicious) network activity, whose generated datapoints are (ii) automatically labeled at the granular level, while also enabling (iii) control of the network conditions (e.g., to simulate resource starvation). As such, ConCap represents a solid foundation to address the “data problem” that affects NIDS research. We will demonstrate this.
@@
To develop ConCap, we assembled open-source technologies, such as container-related frameworks, networking utilities, and network flows (NetFlow [2, 19]) extractors. Our design facilitates experimental reproducibility: to (automatically) generate labeled NetFlows pertaining to one (or more) attacks, ConCap only requires the researcher to define a “scenario,” i.e., a file describing the activities of the “attacker” and “target” host(s), as well as the characteristics of the overarching network channel. In this way, other researchers can replicate the same experiments just by running the same “scenario” in their own version of ConCap, thereby removing the need to share data—which can be quite big in size (e.g., the CICIDS18 [9] dataset is larger than 400GB). Such a property intrinsically fosters scientific reproducibility—which is lacking even in top-tier security venues [20, 2].
@@
We empirically validate the “realisticness” of ConCap’s generated data. Through extensive experiments wherein we compare the network traffic produced by a real-world network with that generated by ConCap, we find that there are negligible differences (at both the packet- and NetFlow-level) between these two setups. We deeply investigate such differences, and confirm that they are due to the well-known “unpredictable” behavior of modern networks [21, 2]. Inspired by such a finding, we go one step further and examine if the data generated by ConCap is deterministic—an aspect that received limited attention from related research (e.g., [22]). Our analysis, entailing 100 executions of the same set of four distinct network activities, reveals that ConCap’s output cannot be claimed to be fully deterministic—as is the case for real-world networks, too. However, through other experiments on well-known benchmarks and on a real “smart-home” network, we show that, ML-wise, the data generated by ConCap is functionally equivalent to that of other network setups.
@@
  • Empirically demonstrate (§V) that ConCap generates network-traffic data resembling that of real-world networks, and prove that experiments run via ConCap are not fully deterministic—a behavior expected by real-world networks.
@@
We also discuss our findings (§VII) and compare ConCap with related works (§VIII). We release all our resources [23].
@@ II Related Work and Motivation
We summarize the domain of NIDS (§II-A) and we outline the challenges of acquiring data for assessing NIDS (§II-B), representing the root of the problem tackled by our paper. Then, we motivate our technical (§II-C) contribution, which specifically focuses on NIDS analysing network flows (NetFlows [24]) due to the widespread usage of this datatype [2, 8, 25].
@@ II-A Network Intrusion Detection Systems (NIDS)
Modern networks are constantly under attack [26, 27], and NIDS represent the first line of defense against the ever-evolving cyberthreats [28]. The main goal of an NIDS [^1] Borrowing from [2], we use “NIDS” in a broad sense, including also SIEM [29] or EDR [30] (which can be seen as extensions of NIDS). is to detect any given threat (e.g., a botnet-infected host, a remote DDoS attack, or an attacker who acquired access to an internal host) as early as possible so that proper mitigations can be enacted to mitigate the damage of an offensive campaign [31]. Fig. 1 shows a schematic of an NIDS deployment scenario.
@@
The advent of data-driven technologies, such as AI/ML (which support both signature- and anomaly-based methods [2]), has been adopted by NIDS developers [32]. Yet, even state-of-the-art NIDSs struggle with the sheer amount of attacks that target current organizations [26]. Intriguingly, even though practitioners do appreciate the analytical capabilities of ML [33], modern security operation centers are overwhelmed with false alarms [34, 35]. Put simply, despite being an instrumental security tool, there is a constant need to improve NIDS—creating a fertile ground for research.
Since the seminal work by Denning [36], thousands of scientific papers have sought to propose (e.g., [37, 38, 39, 40]), enhance (e.g., [41, 42]), or assess (e.g., [16, 43, 44]) NIDS. Despite all such efforts, a recent SoK [2] revealed that there is skepticism among industry experts with regard to the results claimed in research. Such doubts are well-founded: various recent papers (e.g., [6, 7, 8]) identified critical issues that are becoming endemic in research. The root cause of most such critiques is the poor quality of the data used to test these systems—which is a problem that has been known to affect this research domain since at least 2010 [21].
@@ II-B Research Challenge: Obtaining Data for NIDS
Any security tool must be tested before its deployment, and such tests require data. In the context of NIDS, such an evaluation should be carried out on data that is representative of the environment in which the NIDS is meant to be deployed [2]. Unfortunately, carrying out the abovementioned operations is tough to accomplish from the perspective of an (academic) researcher. Let us outline the options that enable one to collect network-related data for a security assessment of a NIDS, explaining their challenges.
  • Real-world capture from the deployment environment. This is ideally the best option since it guarantees that the data resembles the one that will be generated by the monitored network. However, such an option may not be available: the researcher may not have access to such data due to privacy reasons, and infecting/attacking physical devices may not be acceptable in some organisations (even if for research).
  • Synthetic capture from a custom environment. This is a sensible alternative: by creating an ad-hoc network (e.g., via virtual machines [9]), a researcher has plenty of freedom to collect and generate any sort of data. However, the benign data may not be representative of the deployment environment. Moreover, even in such a setup, it is currently challenging to precisely distinguish benign from malicious data points: as shown in [6, 8], there is a risk of mistakenly “label” benign samples as malicious. Such errors can skew the results of the final assessment [45].
  • Reliance on benchmark datasets. The last option is to use publicly available data, e.g., generated by other researchers in their own environments (either from the real-world, or synthetically). This is a convenient option: it is exempt from privacy concerns and requires minimal technical expertise (since no simulation occurs). However, the validity of the corresponding evaluation will depend on whether the benchmark is a meaningful representation of the network wherein the NIDS is meant to be deployed.
For real-world deployments, it is paramount to evaluate the NIDS in the (real) network to be monitored by the NIDS: as highlighted by Sommer and Paxson [21], networks present immense variability. Hence, even if any given NIDS is shown to “work well” on data from a custom network, it is questionable whether the same NIDS works well also in other networks. This is known as the “generalisability” [46, 47] (or “transferability” [48]) problem of NIDS, which prevents the creation of plug-and-play NIDS (confirmed by practitioners [18]). However, a research paper needs not to aim for real-world deployment to provide a significant contribution to the state of the art [2]. Our work is rooted in this truth: we focus on improving future research on NIDS—which not necessarily requires assessments on “real” networks to be valuable.
@@ II-C Practical Generation of Network Data
  • The researcher can create a simulated/virtual network, but doing so requires dealing with the labeling issues (for the malicious traffic), and is (likely) limited to small-scale evaluations due to the impossibility of recreating a large network environment [22].
  • The researcher can exclusively rely on benchmark datasets, but this would limit the evaluation to the data within the benchmark dataset (e.g., even by manipulating the benchmark’s datapoints, there is a risk of breaking domain constraints [49]), preventing exploration of new threats.
  • To overcome the abovementioned limitations, the researcher can do a mix of the above [15, 16]: they can use “benign” data collected from a public dataset (potentially validated by prior work [6]), and then generate “malicious” data to use alongside the benign data to carry out a proper evaluation.
We argue that the third option is the most enticing one. As of 2025, there are various publicly available datasets captured in large networks (see [11] for a list) whose data can be treated as “benign” for experimental purposes. [^2] Such “benign” data can also come from the specific (real) network in which the NIDS is to be deployed—but, as we argued, this is not necessary (although it would increase the soundness of an evaluation). Hence, if one could generate “malicious” data so that (i) it is “correctly labeled,” and (ii) it resembles the data generated by the host of the “benign” network, then one would be able to test a NIDS against a wide array of cyber threats—including new ones.
Research Goal. We seek to develop an open-source tool that enables the automatic generation (and labeling) of network traffic data that resembles a real network.
@@ III-A Objectives
To align these objectives with our overarching research goal (see §II-C), the entirety of ConCap’s source code (which must leverage open-source libraries/frameworks) is publicly released [23]; moreover, the authors will make an effort to ensure long-term maintenance of ConCap’s code repository.
@@ III-B Core Principles
ConCap is rooted on the concept of “scenario”, which serves as a customizable blueprint for an entire experiment. When defining a scenario, developers can specify: (i) the activities carried out by the involved hosts; (ii) the conditions of the overarching network environment; (iii) the fine-grained label to assign to the generated network traffic metadata. We provide in Listing 1 an exemplary “scenario” file (which we will describe in detail in §IV-B), expressed through the YAML notation to ensure broad coverage and understandability.

### v2

# ConCap: Practical Network Traffic Generation for (ML- and) Flow-based Intrusion Detection Systems

@@ I Introduction
There is an anomaly in research on Network Intrusion Detection Systems (NIDS). On the one hand, a deluge of papers continues to expand the boundaries of our knowledge, as shown by khraisat2019survey ; apruzzese2023sok ; ahmad2021network . On the other hand, a growing number of recent efforts highlight some “pitfalls” that undermine the foundations of prior research (e.g., arp2022and ; engelen2021troubleshooting ; liu2022error ; catillo2023machine ; flood2024bad ).
@@
To portray this contrast, we can look at the panorama of open-source datasets used in NIDS-related research, some of which count thousands of citations according to Google Scholar flood2024bad . As a concrete example, consider the paper presenting the CICIDS17 dataset sharafaldin2018toward : published in 2018, this paper had 1600 citations in Q3 2022, which increased to 4200 in Q4 2024—indicating a substantial growth in NIDS research. Yet, in 2021, Engelen et al. engelen2021troubleshooting pinpointed glaring issues in CICIDS17—particularly in terms of ground-truth labeling of attack samples. The flaws of CICIDS17 (and also of its successor, CICIDS18) have been “fixed” in 2022 liu2022error . However, the EuroS&P’24 Best Paper Award flood2024bad revealed that most datasets used by prior research have “bad design smells.”
@@
Simply put, the stark reality is that (i) data is required to support a paper’s claims, but (ii) high-quality data is hard to come by in the NIDS context—especially from the viewpoint of an academic researcher apruzzese2023sok ; flood2024bad ; cordero2021generating . To aggravate this problem, modern NIDS increasingly rely on data-driven techniques, such as machine learning (ML). Therefore, labeled data is necessary to properly evaluate the pros and cons of state-of-the-art NIDS apruzzese2023sok . Finally, despite the benefits that open-source datasets (under the assumption that they are correctly labeled) can provide to a researcher bonninghausen2024introducing , exclusive reliance on such “static benchmarks” prevents one from generating new data. This impairs the assessment of existing NIDS against recent, and more sophisticated threats—such as multi-step attacks navarro2018systematic ; mitreattack , or attacks exploiting recently-discovered vulnerabilities cve . We aim to rectify such a “data problem.”
@@
Our major technical contribution is ConCap, an open-source system to generate (and label) network-traffic data mimicking that of a real-world network. ConCap is particularly suited to generate network data pertaining to malicious activities. Such data can then be used alongside “benign” data taken from the network environment that the NIDS is meant to protect—which is a well-founded assumption arnaldo2019holy ; apruzzese2022cross ; cordero2021generating . Practically, such environment can be: (a) that of a benchmark dataset bonninghausen2024introducing , or of (b) an ad-hoc network testbed for experimental research gomez2023survey , or even that of (c) a real-world network apruzzese2022role . We design ConCap so that it is (i) flexible enough to enable reproduction of any (allegedly malicious) network activity, whose generated datapoints are (ii) automatically labeled at the granular level, while also enabling (iii) control of the network conditions (e.g., to simulate resource starvation). As such, ConCap represents a solid foundation to address the “data problem” that affects NIDS research. We will demonstrate this.
@@
To develop ConCap, we assembled open-source technologies, such as container-related frameworks, networking utilities, and network flows (NetFlow apruzzese2023sok ; verkerken2025rustiflow ) extractors. Our design facilitates experimental reproducibility: to (automatically) generate labeled NetFlows pertaining to one (or more) attacks, ConCap only requires the researcher to define a “scenario,” i.e., a file describing the activities of the “attacker” and “target” host(s), as well as the characteristics of the overarching network channel. In this way, other researchers can replicate the same experiments just by running the same “scenario” in their own version of ConCap, thereby removing the need to share data—which can be quite big in size (e.g., the CICIDS18 sharafaldin2018toward dataset is larger than 400GB). Such a property intrinsically fosters scientific reproducibility—which is lacking even in top-tier security venues olszewski2023get ; apruzzese2023sok .
@@
We empirically validate the realism of ConCap’s generated data. Through extensive experiments wherein we compare the network traffic produced by a physical bare-metal environment with that generated by ConCap, we find that there are negligible differences (at both the packet- and NetFlow-level) between these two setups. We deeply investigate such differences, and confirm that they are due to the well-known “unpredictable” behavior of modern networks sommer2010outside ; apruzzese2023sok . Inspired by such a finding, we go one step further and examine if the data generated by ConCap is deterministic—an aspect that received limited attention from related research (e.g., clausen2019traffic ). Our analysis, entailing 100 executions of the same set of four distinct network activities, reveals that ConCap’s output cannot be claimed to be fully deterministic—as is the case for real-world networks, too. However, through other experiments on well-known benchmarks and on a real “smart-home” network, we show that, ML-wise, the data generated by ConCap is functionally equivalent to that of other network setups.
@@
  • Empirically demonstrate (§V) that ConCap generates network-traffic data resembling that of bare-metal setups, and prove that experiments run via ConCap are not fully deterministic—a behavior expected by real-world networks.
@@
We also discuss our findings (§VII) and compare ConCap with related works (§VIII). We release all our resources repository .
@@ II Related Work and Motivation
We summarize the domain of NIDS (§II-A) and we outline the challenges of acquiring data for assessing NIDS (§II-B), representing the root of the problem tackled by our paper. Then, we motivate our technical (§II-C) contribution, which specifically focuses on NIDS analysing network flows (NetFlows vormayr2020my ) due to the widespread usage of this datatype apruzzese2023sok ; flood2024bad ; dias2020go .
@@ II-A Network Intrusion Detection Systems (NIDS)
Modern networks are constantly under attack enisa2023threat ; cloudflare2024threat , and NIDS represent the first line of defense against the ever-evolving cyberthreats sans2024report . The main goal of an NIDS [^1] Borrowing from apruzzese2023sok , we use “NIDS” in a broad sense, including also SIEM comodo2023idssiem or EDR paloalto2023edrsiem (which can be seen as extensions of NIDS). is to detect any given threat (e.g., a botnet-infected host, a remote DDoS attack, or an attacker who acquired access to an internal host) as early as possible so that proper mitigations can be enacted to mitigate the damage of an offensive campaign bertino2021ai . Fig. 1 shows a schematic of an NIDS deployment scenario.
@@
The advent of data-driven technologies, such as AI/ML (which support both signature- and anomaly-based methods apruzzese2023sok ), has been adopted by NIDS developers kshetri2021economics . Yet, even state-of-the-art NIDSs struggle with the sheer amount of attacks that target current organizations enisa2023threat . Intriguingly, even though practitioners do appreciate the analytical capabilities of ML mink2023everybody , modern security operation centers are overwhelmed with false alarms alahmadi202299 ; vermeer2023alert . Put simply, despite being an instrumental security tool, there is a constant need to improve NIDS—creating a fertile ground for research.
Since the seminal work by Denning denning1987intrusion , thousands of scientific papers have sought to propose (e.g., mirsky2018kitsune ; barradas2021flowlens ; piskozub2021malphase ; fu2021realtime ), enhance (e.g., sommer2003enhancing ; araujo2019improving ), or assess (e.g., apruzzese2022cross ; jacobs2022ai ; corona2013adversarial ) NIDS. Despite all such efforts, a recent SoK apruzzese2023sok revealed that there is skepticism among industry experts with regard to the results claimed in research. Such doubts are well-founded: various recent papers (e.g., liu2022error ; catillo2023machine ; flood2024bad ) identified critical issues that are becoming endemic in research. The root cause of most such critiques is the poor quality of the data used to test these systems—which is a problem that has been known to affect this research domain since at least 2010 sommer2010outside .
@@ II-B Research Challenge: Obtaining Data for NIDS
Any security tool must be tested before its deployment. Such tests require data. For NIDS, the evaluation should be carried out on data that is representative of the environment in which the NIDS is meant to be deployed apruzzese2023sok . Unfortunately, carrying out the abovementioned operations is tough from the perspective of an (academic) researcher. Let us outline the options that enable one to collect network-related data for a security assessment of a NIDS, explaining their challenges.
  • Real-world capture from the deployment environment. This is ideally the best option since it guarantees that the data resembles the one that will be generated by the monitored network. However, such an option may not be available: the researcher may not have access to such data due to privacy reasons, and infecting/attacking physical devices may not be acceptable in some organisations (even if for research wangr2025rr ).
  • Synthetic capture from a custom environment. This is a sensible alternative: by creating an ad-hoc network (e.g., via virtual machines sharafaldin2018toward ), a researcher has plenty of freedom to collect and generate any sort of data. However, the benign data may not be representative of the deployment environment. Moreover, even in such a setup, it is currently challenging to precisely distinguish benign from malicious data points: as shown in liu2022error ; flood2024bad , there is a risk of mistakenly “label” benign samples as malicious. Such errors can skew the results of the final assessment krauss2024verify .
  • Reliance on benchmark datasets. The last option is to use publicly available data, e.g., generated by other researchers in their own environments (either from physical or virtualized setups). This is a convenient option: it is exempt from privacy concerns and requires minimal technical expertise (since no simulation occurs). However, the validity of the corresponding evaluation will depend on whether the benchmark is a meaningful representation of the network wherein the NIDS is meant to be deployed.
For real-world deployments, it is paramount to evaluate the NIDS in the (real) network to be monitored by the NIDS: as highlighted by Sommer and Paxson sommer2010outside , networks present immense variability. Hence, even if any given NIDS is shown to “work well” on data from a custom network, it is questionable whether the same NIDS works well also in other networks. This is known as the “generalisability” verkerken2022towards ; sarhan2022evaluating (or “transferability” catillo2022transferability ) problem of NIDS, which prevents the creation of plug-and-play NIDS (confirmed by practitioners apruzzese2022role ). However, a research paper needs not to aim for real-world deployment to provide a significant contribution to the state of the art apruzzese2023sok . Our work is rooted in this truth: we focus on improving future research on NIDS—which not necessarily requires assessments on “real” networks to be valuable.
@@ II-C Practical Generation of Network Data
  • The researcher can create a simulated/virtual network, but doing so requires dealing with the labeling issues (for the malicious traffic), and is (likely) limited to small-scale evaluations due to the impossibility of recreating a large network environment clausen2019traffic .
  • The researcher can exclusively rely on benchmark datasets, but this would limit the evaluation to the data within the benchmark dataset (e.g., even by manipulating the benchmark’s datapoints, there is a risk of breaking domain constraints sheatsley2021robustness ), preventing exploration of new threats.
  • To overcome the abovementioned limitations, the researcher can do a mix of the above arnaldo2019holy ; apruzzese2022cross : they can use “benign” data collected from a public dataset (potentially validated by prior work liu2022error ), and then generate “malicious” data to use alongside the benign data to carry out a proper evaluation.
We argue that the third option is the most enticing one. As of 2025, there are various publicly available datasets captured in large networks (see bonninghausen2024introducing for a list) whose data can be treated as “benign” for experimental purposes. [^2] Such “benign” data can also come from the specific (real) network in which the NIDS is to be deployed—but, as we argued, this is not necessary (although it would increase the soundness of an evaluation). Hence, if one could generate “malicious” data so that (i) it is “correctly labeled,” and (ii) it resembles the data generated by the host of the “benign” network, then one would be able to test a NIDS against a wide array of cyber threats—including new ones.
Research Goal. We seek to develop an open-source tool that enables the automatic generation (and labeling) of network traffic data that resembles that of a real network.
@@ III-A Objectives
To align these objectives with our overarching research goal (see §II-C), the entirety of ConCap’s source code (which must leverage open-source libraries/frameworks) is publicly released repository ; moreover, the authors will make an effort to ensure long-term maintenance of ConCap’s code repository.
@@ III-B Core Principles
ConCap is rooted on the concept of “scenario”, which serves as a customizable blueprint for an entire experiment. When defining a scenario, developers can specify: (i) the activities carried out by the involved hosts; (ii) the conditions of the overarching network environment; (iii) the fine-grained label to assign to the generated network traffic metadata.